Consider an LC circuit with L = 13 mH and C = 2352 pF. At what frequency is the reactance of the inductor equal to the reactance of the capacitor?

=_____Hz

What is the resonant frequency of the circuit?

=_____Hz
Answer #1

reactance of inductor XL = 2pifL

reacatnce of capacitor Xc = 1/(2pifC)

so XL = Xc

2pifL = 1/2pifC

f^2 = 1/(4pi^2 LC)

f = 1/(2pi sqrt(LC)

f = 1/(2*3.14*(sqrt(0.013* 0.002352)

f = 28.79 KHz
-------------------------------------------
resonant frequency f = 28.79 KHz